Find the domain of the rational function:
(y+5)/(y^2-5y)

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

You cannot divide by zero. So y^2 - 5y can't be zero.

If it were, then

y^2 - 5y = 0

y(y - 5) = 0

y = 0 or y - 5 = 0

y = 0 or y = 5

Those last two values make the denominator zero. So we must kick them out of the domain.

So the domain is the set of all real numbers except y can't be 0 or y can't be 5.

In set-builder notation, the domain is

In interval notation, the domain is

Answer by stanbon(75887)

Find the domain of the rational function:
(y+5)/(y^2-5y)
----
The denominator cannot be zero.
Solve:
y^2-5y = 0
Factor:
y(y-5) = 0
y = 0 or y = 5
----
Ans: Domain is All Real Numbers except y = 0 and y = -5.
=============
Cheers,
Stan H.
=================